# Finite Difference Solution Ansatz approach in Least-Squares Monte Carlo

### v8

@@ 2.1 Model Settings
with $\operatorname{\mathbb{E}}[dB^{S}(t)dB^{\nu}(t)]=\rho^{S,\nu}dt$, where $q$ is the constant continuous dividend yield, $\theta$ is the long variance, $\kappa$ is the mean reverting rate, $\xi$ is the volatility of volatility and $\rho^{S,\nu}$ is the correlation between spot and variance. For the same reason as previous subsection, we assume $S(0)=1$ for simplicity. And the initial variance is given by $\nu(0)$. There is a well known result on the expectation value of $\nu(t)$ over time: $\operatorname{\mathbb{E}}[\nu(t)]=\nu(0)e^{-\kappa t}+\theta(1-e^{-\kappa t})$.
@@ 2.2.1 Problem Formulation
As a general framework relevant for derivative valuation as well as regulatory CVA computations, we assume that throughout a given terminal expiry $T$ with the riskless interest rate $r$ defined previously, there exists a complete probability space ($\Omega,\operatorname{\mathcal{F}},(\operatorname{\mathcal{F}}_{t})_{0\leq t\leq T},\operatorname*{\mathbb{Q}}$), accommodating the model settings in Sec.2.1, under risk-neutral measure $\operatorname*{\mathbb{Q}}$. Also, let $L^{2}(\Omega,\operatorname{\mathcal{F}},(\operatorname{\mathcal{F}}_{t})_{0\leq t\leq T},\operatorname*{\mathbb{Q}})$, hereinafter simplified as $L^{2}$, denote the space of square integrable functions. As problem formulation, from $t=0$ and final expiry $T$ of an derivative contract, we define a set of discretized time points as $\pi(t):=\{t_{k}\geq t|t_{k}=k\Delta T,1\leq k\leq M=\frac{T}{\Delta T}\}$ with a time interval $\Delta T$. $\pi(t)$ represents either the exercise times for American-style option, or discrete exposure monitoring times in CVA calculation. Within the complete probability space $\Omega$, there exists a discrete filtration $\{\operatorname{\mathcal{F}}(t_{k})|1\leq k\leq M\}$ on $\pi(0)$. For notational clarity, we use $XYZ(\omega,t)$ to denote a stochastic variable under $\operatorname{\mathcal{F}}(t)$ with $\omega\in\Omega$. In this spirit, the state variables are denoted by $X(\omega,t_{k})$, which is $\operatorname{\mathcal{F}}(t_{k})$-measurable.
@@ 2.2.4 Finite difference approach in PDE
As opposed to LSM, alternative numerical approach to solve the American-style option pricing is the finite difference (FD) method. The elegance of this method lies in the fact that it marches from maturity to present, which naturally fits into the requirement of backward recursive optimization in stopping time problems. We restrict the discussion within 1D for maximum efficiency and accuracy. The local volatility setting as describe in Sec.2.1 with $d=1$ is solvable under 1D FD PDE, under which the option price $V^{FD}(S,t)$ as function of underlying spot $S$ (omitting the subscript $i$ for the mono underlying case) and time $t$ satisfies:
@@ 2.3.2 FD Solution: Control Variate
From the point on, for notational convenience on random variables, we suppress the spatial dependence and condense the notation by replacing arguments depending on time by subscripts and further replace $t_{k}$ with simply $k$ if applicable. For example, $X_{k}\equiv X(\omega,t_{k})$, $\vec{\phi}^{P}_{R,k}\equiv\vec{\phi}^{P}_{R}(X_{k})$, $\vec{\phi}^{P^{\text{FD}}}_{k}\equiv\vec{\phi}^{P^{\text{FD}}}(X_{k})$, $V_{k+1}\equiv V(\omega,\tau(t_{k+1}))$, $Z_{k}\equiv Z(\omega,t_{k})$, $F_{k}\equiv F(\omega,t_{k})$ and $F^{P}_{k}\equiv F^{P}(\omega,t_{k})$.
@@
$\begin{split}\beta_{k}^{\text{FD}}(Y_{t})&:=\frac{\operatorname{\mathbb{E}}[\operatorname{\mathrm{Cov}}(Y_{t},\hat{V}_{k+1}^{1D}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]}{\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(\hat{V}_{k+1}^{1D}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]},\\ \alpha_{k}^{\text{FD}}(Y_{t})&:=\operatorname{\mathbb{E}}(Y_{t}|\operatorname{\mathcal{F}}_{0})-\beta_{k}^{\text{FD}}(Y_{t})e^{-r\Delta T}\operatorname{\mathbb{E}}(\hat{V}_{k+1}^{1D}|\operatorname{\mathcal{F}}_{0}),\end{split}$ (2.23)
@@
because $\beta_{k}^{\text{FD}}(F_{k})=\beta_{k}^{\text{FD}}(e^{-r\Delta T}V_{k+1})$ and $\alpha_{k}^{\text{FD}}(F_{k})=\alpha_{k}^{\text{FD}}(e^{-r\Delta T}V_{k+1})$ given the $\operatorname{\mathcal{F}}_{t_{k}}$ filtration.
We also define the annihilator under $\vec{\phi}^{P^{\text{FD}}}(x)$ as $M_{k}^{\text{FD}}:=1-P_{k}^{\text{FD}}$. Both of $P_{k}^{\text{FD}}$ and $M_{k}^{\text{FD}}$ are idempotent.
@@
$P_{k}^{\text{FD-LSM}}$ on the option price can be decomposed exactly into $P_{k}^{\text{FD}}$ (defined in above) and $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$: $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$ utilizes the same basis functions as $P_{k}^{\text{LSM}}$, but the regression coefficients are obtained as the “net” effect of $\vec{\phi}_{R-1}^{P^{\text{LSM}}}(x)$ on $V_{k+1}$ with the effect of $\vec{\phi}^{P^{\text{FD}}}(x)$ annihilated from the remaining system.
@@
Here we are trying to partition a $L^{2}$ projection into 2 sets of explanatory variables, $\vec{\phi}^{P^{\text{FD}}}(x)$ and $\vec{\phi}_{R-1}^{P^{\text{LSM}}}(x)$. Based on the so-called Frisch-Waugh-Lovell (FWL) theorem  [^5] This theorem is well-recognized in the field of Econometrics, as a powerful tool for understanding and simplifying regression analysis. Its proof can be found in Section 2.4 of Davidson et al., (2004)., instead of regressing on all variables as a whole, one can perform two subsequent regressions on individual set: while the regression on $\vec{\phi}^{P^{\text{FD}}}(x)$ is done as usual, the regression coefficients of $\vec{\phi}_{R-1}^{P^{\text{LSM}}}(x)$ can be obtained equivalently by first “partialling out” the effects of $\vec{\phi}^{P^{\text{FD}}}(x)$ from both $V_{k+1}$ and $\vec{\phi}_{R-1}^{P^{\text{LSM}}}(x)$, and then running the regression of the partialled-out dependent variable, denoted by $V^{\perp\text{FD}}_{k+1}$, on the partialled-out regressors, denoted by $\vec{\phi}^{P^{\text{LSM}\perp\text{FD}}}_{R-1}(x)$.
In mathematical terms, $P_{k}^{\text{FD-LSM}}$ can be decomposed into $P_{k}^{\text{FD}}$ and $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$, where the latter is defined as
@@
Here $\mathcal{L}$ adopts the same functional form as Eqn.2.5 but operates on below arguments by annihilating the effect of $\vec{\phi}^{P^{\text{FD}}}(x)$
@@
Due to the idempotent property of $M_{k}^{\text{FD}}$, we have $\vec{\beta}^{P^{\text{LSM}\perp\text{FD}}}_{k}(e^{-r\Delta T}V_{k+1})=\vec{\beta}^{P^{\text{LSM}\perp\text{FD}}}_{k}(e^{-r\Delta T}V_{k+1}^{\perp\text{FD}})$ and $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}V_{k+1}]=P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}V^{\perp\text{FD}}_{k+1}]$. Thus, we arrive at
$\begin{split}P_{k}^{\text{FD-LSM}}[e^{-r\Delta T}V_{k+1}]&=P_{k}^{\text{FD}}[e^{-r\Delta T}V_{k+1}]+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}V_{k+1}]\\ &=P_{k}^{\text{FD}}[e^{-r\Delta T}V_{k+1}]+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}V_{k+1}^{\perp\text{FD}}].\end{split}$ (2.28)
@@
First of all, based on Lemma 2, let’s apply FD-LSM to the path option value under 1D process, which gives,
$\begin{split}P_{k}^{\text{FD-LSM}}[e^{-r\Delta T}\hat{V}_{k+1}^{1D}]&=P_{k}^{\text{FD}}[e^{-r\Delta T}\hat{V}_{k+1}^{1D}]+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}\hat{V}_{k+1}^{1D}-P_{k}^{\text{FD}}[e^{-r\Delta T}\hat{V}_{k+1}^{1D}]]\\ &=f_{k}^{\text{FD}}(X_{k})+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}\hat{V}_{k+1}^{1D}-f_{k}^{\text{FD}}(X_{k})].\end{split}$ (2.29)
@@
$\begin{split}P_{k}^{\text{FD-LSM}}[e^{-r\Delta T}V_{k+1}]&=P_{k}^{\text{FD}}[e^{-r\Delta T}V_{k+1}]+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}V_{k+1}-P_{k}^{\text{FD}}[e^{-r\Delta T}V_{k+1}]]\\ &=\beta_{k}^{\text{FD}}(e^{-r\Delta T}V_{k+1})f_{k}^{\text{FD}}(X_{k})+\alpha_{k}^{\text{FD}}(e^{-r\Delta T}V_{k+1})\\ &\quad+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}V_{k+1}-\beta_{k}^{\text{FD}}(e^{-r\Delta T}V_{k+1})f_{k}^{\text{FD}}(X_{k})-\alpha_{k}^{\text{FD}}(e^{-r\Delta T}V_{k+1})]\\ &=\beta_{k}^{\text{FD}}(F_{k})f_{k}^{\text{FD}}(X_{k})+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})f_{k}^{\text{FD}}(X_{k})],\end{split}$ (2.31)
where in the last line, $\beta_{k}^{\text{FD}}(e^{-r\Delta T}V_{k+1})$ has been replaced with $\beta_{k}^{\text{FD}}(F_{k})$ equivalently from Eqn.2.23 that leads to Eqn.2.24, and constant term $\alpha_{k}^{\text{FD}}(F_{k})$ has been absorbed into $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$.
Given Eqn.2.30, we have $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}\hat{V}_{k+1}^{1D}]=P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[f_{k}^{\text{FD}}(X_{k})]$. Thus we arrive at the desired result as a consequence of linearity property of $L^{2}$ projection:
$P_{k}^{\text{FD-LSM}}[e^{-r\Delta T}V_{k+1}]=\beta_{k}^{\text{FD}}(F_{k})f_{k}^{\text{FD}}(X_{k})+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D})],$ (2.32)
which claims that $P_{k}^{\text{FD-LSM}}$ is equivalent to a customized projection on a control-variated “kernel” with the help of the FD ansatz. ∎
Based on Prop.1, proof of convergence on LSM from previous studies (Longstaff and Schwartz, (2001); Stentoft, (2001)) is equally applicable to FD-LSM. Both $P_{k}^{\text{FD-LSM}}$ and $P_{k}^{\text{LSM}}$ are unbiased estimators. However, the variance of projection “kernel” in the former is reduced tremendously in a control variated manner by observing
@@
Considering Eqn.2.7 (re-labeling $P$ to $P^{\text{LSM}}$) and Eqn.2.26 together with above, one can see that the RSS (lower bound) in FD-LSM is also reduced by the same scaling factor as the variance:
$\begin{split}\inf_{\vec{b}\in\mathbb{R}^{R}}\mathcal{L}(e^{-r\Delta T}V_{k+1};\vec{b};\vec{\phi}^{P^{\text{FD-LSM}}}_{R,k})&=\inf_{\vec{b}\in\mathbb{R}^{R-1}}\mathcal{L}(e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D});\vec{b};\vec{\phi}^{P^{\text{LSM}\perp\text{FD}}}_{R-1,k})\\ &=\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(e^{-r\Delta T}V_{k+1}-e^{-r\Delta T}\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]\\ &=[1-\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}_{k+1}^{1D})]\cdot\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(e^{-r\Delta T}V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]\\ &=[1-\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}_{k+1}^{1D})]\cdot\inf_{\vec{b}\in\mathbb{R}^{R}}\mathcal{L}(e^{-r\Delta T}V_{k+1};\vec{b};\vec{\phi}^{P^{\text{LSM}}}_{R,k}).\end{split}$ (2.35)
It means the stronger the correlation between $V_{k+1}$ and $\hat{V}_{k+1}^{1D}$, the more effective the error reduction achieved in the new scheme; as always, such error reduction directly translates into the decrease in $N_{R}$ required to fulfill suitable precision.

### v9

@@ 2.1 Model Settings
with $\operatorname{\mathbb{E}}[dB^{S}(t)dB^{\nu}(t)]=\rho^{S,\nu}dt$, where $q$ is the constant continuous dividend yield, $\theta$ is the long variance, $\kappa$ is the mean reverting rate, $\xi$ is the volatility of volatility and $\rho^{S,\nu}$ is the correlation between spot and variance. For the same reason as previous subsection, we assume $S(0)=1$ for simplicity. Given the initial variance $\nu(0)$, there is a well known result on the expectation value of $\nu(t)$ over time: $\operatorname{\mathbb{E}}[\nu(t)]=\nu(0)e^{-\kappa t}+\theta(1-e^{-\kappa t})$.
@@ 2.2.1 Problem Formulation
As a general framework relevant for derivative valuation as well as regulatory CVA computations, we assume that throughout a given terminal expiry $T$ with the riskless interest rate $r$ defined previously, there exists a complete probability space ($\Omega,\operatorname{\mathcal{F}},(\operatorname{\mathcal{F}}_{t})_{0\leq t\leq T},\operatorname*{\mathbb{Q}}$), accommodating the model settings in Sec.2.1, under risk-neutral measure $\operatorname*{\mathbb{Q}}$. Also, let $L^{2}(\Omega,\operatorname{\mathcal{F}},(\operatorname{\mathcal{F}}_{t})_{0\leq t\leq T},\operatorname*{\mathbb{Q}})$, hereinafter simplified as $L^{2}$, denote the space of square integrable functions. As problem formulation, from $t=0$ and final expiry $T$ of an derivative contract, we define a set of discretized time points as $\pi(t):=\{t_{k}\geq t|t_{k}=k\Delta T,1\leq k\leq M=\frac{T}{\Delta T}\}$ with a time interval $\Delta T$. $\pi(t)$ represents either the exercise times for American-style option, or discrete exposure monitoring times in CVA calculation. Within the complete probability space $\Omega$, there exists a discrete filtration $\{\operatorname{\mathcal{F}}_{t_{k}}|1\leq k\leq M\}$ on $\pi(0)$. For notational clarity, we use $XYZ(\omega,t)$ to denote a stochastic variable under $\operatorname{\mathcal{F}}_{t}$ with $\omega\in\Omega$. In this spirit, the state variables are denoted by $X(\omega,t_{k})$, which is $\operatorname{\mathcal{F}}_{t_{k}}$-measurable.
@@ 2.2.4 Finite difference approach in PDE
As opposed to LSM, alternative numerical approach to solve the American-style option pricing is the finite difference (FD) method. The elegance of this method lies in the fact that it marches from maturity to present, which naturally fits into the requirement of backward recursive optimization in stopping time problems. We restrict the discussion within 1D for maximum efficiency. The local volatility setting as describe in Sec.2.1 with $d=1$ is solvable under 1D FD PDE, under which the option price $V^{FD}(S,t)$ as function of underlying spot $S$ (omitting the subscript $i$ for the mono underlying case) and time $t$ satisfies:
@@ 2.3.1 Limiting Case
Before further elaborating on FD-LSM, we introduce simplified notation for random variables to enhance clarity and conciseness from now on. Specifically, we suppress the spatial dependence and condense the notation by replacing arguments depending on time by subscripts, and then further replace $t_{k}$ with simply $k$ whenever applicable. For example, $X_{k}\equiv X(\omega,t_{k})$, $\vec{\phi}^{P}_{R,k}\equiv\vec{\phi}^{P}_{R}(X_{k})$, $\vec{\phi}^{P^{\text{FD}}}_{k}\equiv\vec{\phi}^{P^{\text{FD}}}(X_{k})$ (to be defined in next section), $V_{k+1}\equiv V(\omega,\tau(t_{k+1}))$, $Z_{k}\equiv Z(\omega,t_{k})$, $F_{k}\equiv F(\omega,t_{k})$ and $F^{P}_{k}\equiv F^{P}(\omega,t_{k})$.
@@ 2.3.2 FD Solution: Control Variate
$\begin{split}\beta_{k}^{\text{FD}}(Y_{t})&:=\frac{\operatorname{\mathbb{E}}[\operatorname{\mathrm{Cov}}(e^{r\Delta T}Y_{t},\hat{V}_{k+1}^{1D}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]}{\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(\hat{V}_{k+1}^{1D}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]},\\ \alpha_{k}^{\text{FD}}(Y_{t})&:=\operatorname{\mathbb{E}}(Y_{t}|\operatorname{\mathcal{F}}_{0})-\beta_{k}^{\text{FD}}(Y_{t})e^{-r\Delta T}\operatorname{\mathbb{E}}(\hat{V}_{k+1}^{1D}|\operatorname{\mathcal{F}}_{0}),\end{split}$ (2.23)
@@
because $\beta_{k}^{\text{FD}}(F_{k})=\beta_{k}^{\text{FD}}(e^{-r\Delta T}V_{k+1})$ and $\alpha_{k}^{\text{FD}}(F_{k})=\alpha_{k}^{\text{FD}}(e^{-r\Delta T}V_{k+1})$ given the identity $F_{k}=\operatorname{\mathbb{E}}(e^{-r\Delta T}V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}})$.
We also define the annihilator under $\vec{\phi}^{P^{\text{FD}}}(x)$ as $M_{k}^{\text{FD}}:=1-P_{k}^{\text{FD}}$. Both of $P_{k}^{\text{FD}}$ and $M_{k}^{\text{FD}}$ are idempotent, i.e., $P_{k}^{\text{FD}}[P_{k}^{\text{FD}}[\cdot]]=P_{k}^{\text{FD}}[\cdot]$ and $M_{k}^{\text{FD}}[M_{k}^{\text{FD}}[\cdot]]=M_{k}^{\text{FD}}[\cdot]$.
@@
$P_{k}^{\text{FD-LSM}}$ on the option price can be decomposed exactly into $P_{k}^{\text{FD},\text{FD}\perp\text{LSM}}$ and $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$: $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$ employs the same basis functions as $P_{k}^{\text{LSM}}$, but the regression coefficients are obtained as the “net” effect of $\vec{\phi}_{R-1}^{P^{\text{LSM}}}(x)$ on $V_{k+1}$ with the effect of $\vec{\phi}^{P^{\text{FD}}}(x)$ annihilated from the remaining system; $P_{k}^{\text{FD},\text{FD}\perp\text{LSM}}$ is defined analogously, but with the roles of $\vec{\phi}^{P^{\text{FD}}}(x)$ and $\vec{\phi}_{R-1}^{P^{\text{LSM}}}(x)$ interchanged.
@@
Here, we decompose an $L^{2}$ projection using two sets of basis functions, $\vec{\phi}^{P^{\text{FD}}}(x)$ and $\vec{\phi}_{R-1}^{P^{\text{LSM}}}(x)$, effectively implementing a Frisch-Waugh-Lovell (FWL) type decomposition into and $P_{k}^{\text{FD},\text{FD}\perp\text{LSM}}$ and $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$. We will focus on a detailed examination of $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$, noting that $P_{k}^{\text{FD},\text{FD}\perp\text{LSM}}$ can be analyzed symmetrically by swapping the roles of the FD and LSM basis functions.
Based on the FWL theorem [^5] This theorem is well-recognized in the field of Econometrics, as a powerful tool for understanding and simplifying regression analysis. Its proof is provided in Section 2.4 of Davidson et al., (2004), or more recently in Lovell, (2008)., the mathematical definition of $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$ is
@@
Here $\mathcal{L}$ adopts the same functional form as Eqn.2.5 but operates on below arguments by partialling out the effect of $\vec{\phi}^{P^{\text{FD}}}(X_{k})$
@@
Given the symmetrical nature of $\vec{\phi}^{P^{\text{FD}}}(x)$ and $\vec{\phi}_{R-1}^{P^{\text{LSM}}}(x)$ with respect to the applicability of the FWL theorem, $P_{k}^{\text{FD},\text{FD}\perp\text{LSM}}[e^{-r\Delta T}V_{k+1}]$ is defined as a modified $P_{k}^{\text{FD}}[e^{-r\Delta T}V_{k+1}]$, controlling for $\vec{\phi}_{R-1}^{P^{\text{LSM}}}(X_{k})$:
$\begin{split}P_{k}^{\text{FD},\text{FD}\perp\text{LSM}}[e^{-r\Delta T}V_{k+1}]&:=\beta_{k}^{\text{FD}\perp\text{LSM}}(e^{-r\Delta T}V_{k+1})f_{k}^{\text{FD}}(X_{k})+\alpha_{k}^{\text{FD}\perp\text{LSM}}(e^{-r\Delta T}V_{k+1}),\\ \beta_{k}^{\text{FD}\perp\text{LSM}}(e^{-r\Delta T}V_{k+1})&:=\frac{\operatorname{\mathbb{E}}[\operatorname{\mathrm{Cov}}(V_{k+1}^{\perp\text{LSM}},\hat{V}_{k+1}^{1D\perp\text{LSM}}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]}{\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(\hat{V}_{k+1}^{1D\perp\text{LSM}}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]},\\ \alpha_{k}^{\text{FD}\perp\text{LSM}}(e^{-r\Delta T}V_{k+1})&:=e^{-r\Delta T}(\operatorname{\mathbb{E}}(V_{k+1}^{\perp\text{LSM}}|\operatorname{\mathcal{F}}_{0})-\beta_{k}^{\text{FD}\perp\text{LSM}}(e^{-r\Delta T}V_{k+1})\operatorname{\mathbb{E}}(\hat{V}_{k+1}^{1D\perp\text{LSM}}|\operatorname{\mathcal{F}}_{0})),\end{split}$ (2.28)
where $e^{-r\Delta T}Y_{k+1}^{\perp\text{LSM}}$, with $Y_{k+1}$ being either $V_{k+1}$ or $\hat{V}_{k+1}^{1D}$, is defined as:
$\begin{split}e^{-r\Delta T}Y_{k+1}^{\perp\text{LSM}}:&=e^{-r\Delta T}Y_{k+1}-\arg_{\vec{b}\in\mathbb{R}^{R-1}}\min\mathcal{L}(e^{-r\Delta T}Y_{k+1};\vec{b};\vec{\phi}_{R-1}^{P^{\text{LSM}}}(X_{k}))\cdot\vec{\phi}_{R-1}^{P^{\text{LSM}}}(X_{k})\\ &=e^{-r\Delta T}Y_{k+1}-(\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R-1,k}\vec{\phi}^{P^{\text{LSM}}T}_{R-1,k}|\operatorname{\mathcal{F}}_{0}])^{-1}\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R-1,k}e^{-r\Delta T}Y_{k+1}|\operatorname{\mathcal{F}}_{0}]\cdot\vec{\phi}_{R-1}^{P^{\text{LSM}}}(X_{k}).\end{split}$
Combining Eqn.2.25 and 2.28, we manage to arrive at
$\begin{split}P_{k}^{\text{FD-LSM}}[e^{-r\Delta T}V_{k+1}]&=P_{k}^{\text{FD},\text{FD}\perp\text{LSM}}[e^{-r\Delta T}V_{k+1}]+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}V_{k+1}]\\ &=P_{k}^{\text{FD},\text{FD}\perp\text{LSM}}[e^{-r\Delta T}V_{k+1}]+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}V_{k+1}^{\perp\text{FD}}],\end{split}$ (2.29)
where in the last line we have applied $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}V_{k+1}]=P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}V^{\perp\text{FD}}_{k+1}]$, induced by the idempotent property of $M_{k}^{\text{FD}}$, to facilitate further discussion.
@@
First of all, based on Lemma 2, let’s apply FD-LSM in Eqn.2.29 to the path option value under 1D process, which gives,
$\begin{split}P_{k}^{\text{FD-LSM}}[e^{-r\Delta T}\hat{V}_{k+1}^{1D}]&=P_{k}^{\text{FD},\text{FD}\perp\text{LSM}}[e^{-r\Delta T}\hat{V}_{k+1}^{1D}]+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}\hat{V}_{k+1}^{1D}-P_{k}^{\text{FD}}[e^{-r\Delta T}\hat{V}_{k+1}^{1D}]]\\ &=f_{k}^{\text{FD}}(X_{k})+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}\hat{V}_{k+1}^{1D}-f_{k}^{\text{FD}}(X_{k})],\end{split}$ (2.30)
where we have utilized the identities $\beta_{k}^{\text{FD}\perp\text{LSM}}(e^{-r\Delta T}\hat{V}_{k+1}^{1D})=\beta_{k}^{\text{FD}}(e^{-r\Delta T}\hat{V}_{k+1}^{1D})=1$ and $\alpha_{k}^{\text{FD}\perp\text{LSM}}(e^{-r\Delta T}\hat{V}_{k+1}^{1D})=\alpha_{k}^{\text{FD}}(e^{-r\Delta T}\hat{V}_{k+1}^{1D})=0$.
@@
$\begin{split}P_{k}^{\text{FD-LSM}}[e^{-r\Delta T}V_{k+1}]&=P_{k}^{\text{FD},\text{FD}\perp\text{LSM}}[e^{-r\Delta T}V_{k+1}]+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}V_{k+1}-P_{k}^{\text{FD}}[e^{-r\Delta T}V_{k+1}]]\\ &=\beta_{k}^{\text{FD}\perp\text{LSM}}(e^{-r\Delta T}V_{k+1})f_{k}^{\text{FD}}(X_{k})+\alpha_{k}^{\text{FD}\perp\text{LSM}}(e^{-r\Delta T}V_{k+1})\\ &\quad+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}V_{k+1}-\beta_{k}^{\text{FD}}(e^{-r\Delta T}V_{k+1})f_{k}^{\text{FD}}(X_{k})-\alpha_{k}^{\text{FD}}(e^{-r\Delta T}V_{k+1})]\\ &=\beta_{k}^{\text{FD}\perp\text{LSM}}(e^{-r\Delta T}V_{k+1})f_{k}^{\text{FD}}(X_{k})+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})f_{k}^{\text{FD}}(X_{k})],\end{split}$ (2.32)
where in the last line, we have employed the identity $\beta_{k}^{\text{FD}}(e^{-r\Delta T}V_{k+1})=\beta_{k}^{\text{FD}}(F_{k})$ from Eqn.2.24, and the constant term $\alpha_{k}^{\text{FD}\perp\text{LSM}}(\cdot)-\alpha_{k}^{\text{FD}}(\cdot)$ has been absorbed into $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$.
Given Eqn.2.31, we have $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}\hat{V}_{k+1}^{1D}]=P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[f_{k}^{\text{FD}}(X_{k})]$. Thus we arrive at the desired result as a consequence of linearity property of $L^{2}$ projection:
$P_{k}^{\text{FD-LSM}}[e^{-r\Delta T}V_{k+1}]=\beta_{k}^{\text{FD}\perp\text{LSM}}(e^{-r\Delta T}V_{k+1})f_{k}^{\text{FD}}(X_{k})+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D})].$ (2.33)
The following Sec.2.3.3 will demonstrate that the primary source of regression error in this decomposition arises from the second term, equivalent to a customized projection on a control-variated “kernel” with the help of the FD ansatz. ∎
Based on Prop.1, proof of convergence on LSM from previous studies (Longstaff and Schwartz, (2001); Stentoft, (2001)) is equally applicable to FD-LSM. Both $P_{k}^{\text{FD-LSM}}$ and $P_{k}^{\text{LSM}}$ are unbiased estimators. However, the variance of projection “kernel” in the former is reduced tremendously in a control variated manner by observing (with the help of Eqn.2.23)
@@
Comparing the objective function in Eqn.2.26 with Eqn.2.7 (re-labeling $P$ to $P^{\text{LSM}}$), and applying the aforementioned identity, one can see that the scaling factor that reduces the variance in the projection $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$ also proportionally reduces the RSS (lower bound):
$\begin{split}\inf_{\vec{b}\in\mathbb{R}^{R-1}}\mathcal{L}(e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D});\vec{b};\vec{\phi}^{P^{\text{LSM}\perp\text{FD}}}_{R-1,k})&=\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(e^{-r\Delta T}V_{k+1}-e^{-r\Delta T}\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]\\ &=[1-\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}_{k+1}^{1D})]\cdot\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(e^{-r\Delta T}V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]\\ &=[1-\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}_{k+1}^{1D})]\cdot\inf_{\vec{b}\in\mathbb{R}^{R}}\mathcal{L}(e^{-r\Delta T}V_{k+1};\vec{b};\vec{\phi}^{P^{\text{LSM}}}_{R,k}).\end{split}$
It means the stronger the correlation between $V_{k+1}$ and $\hat{V}_{k+1}^{1D}$, the more effective the variance reduction achieved in the new scheme; as always, such reduction directly translates into the decrease in $N_{R}$ required to fulfill suitable precision.